Calculate the number of Li+ ions in 8.12g Li3PO4


1
Expert's answer
2021-10-15T02:58:08-0400

Since lithium orthophosphate contains 3 lithium ions, the amount of lithium ion substance will be equal to:

"n(Li3PO4) = \frac{9.18}{115.79} = 0.06 \ mole"

"n(Li^+) = n(Li3PO4) \times 3 = 0.18 \ mole"